### v3

# HealthQ: Unveiling Questioning Capabilities of LLM Chains in Healthcare Conversations

Abstract
In digital healthcare, large language models (LLMs) have primarily been utilized to enhance question-answering capabilities and improve patient interactions. However, effective patient care necessitates LLM chains that can actively gather information by posing relevant questions. This paper presents HealthQ, a novel framework designed to evaluate the questioning capabilities of LLM healthcare chains. We implemented several LLM chains, including Retrieval-Augmented Generation (RAG), Chain of Thought (CoT), and reflective chains, and introduced an LLM judge to assess the relevance and informativeness of the generated questions. To validate HealthQ, we employed traditional Natural Language Processing (NLP) metrics such as Recall-Oriented Understudy for Gisting Evaluation (ROUGE) and Named Entity Recognition (NER)-based set comparison, and constructed two custom datasets from public medical note datasets, ChatDoctor and MTS-Dialog. Our contributions are threefold: we provide the first comprehensive study on the questioning capabilities of LLMs in healthcare conversations, develop a novel dataset generation pipeline, and propose a detailed evaluation methodology.
keywords:
Artificial Intelligence in Healthcare , Biomedical Informatics , Digital Health , Natural Language Processing , Information Retrieval in Medicine , Medical Conversational AI , Healthcare Informatics
[…]
[UCI]organization=University of California, Irvine, addressline=Department of Computer Science, city=Irvine, postcode=92697, state=CA, country=USA
[…]
[UCI2]organization=University of California, Irvine, addressline=School of Nursing, city=Irvine, postcode=92697, state=CA, country=USA
[…]
[Columbia]organization=Columbia University, addressline=Department of Statistics, city=New York, postcode=10027, state=NY, country=USA
[…]
[WUSTL]organization=Washington University in St. Louis, addressline=Department of Computer Science and Engineering, city=St. Louis, postcode=63130, state=MO, country=USA
1 Introduction
The integration of Artificial Intelligence (AI) in healthcare has significantly transformed the landscape of personal healthcare, enabling advancements in diagnostic accuracy, treatment personalization, and patient management [1, 2, 3, 4]. Central to these advancements are Large Language Models (LLMs), which have shown outstanding abilities in comprehending and producing human-like text. These capabilities have been harnessed to facilitate various healthcare interventions, from virtual health assistants to automated diagnostic systems [5]. Current research in digital healthcare has extensively explored the use of LLMs to enhance question-answering (QA) capabilities. QA systems are crucial as they provide patients with accurate and timely responses to their medical inquiries, thereby improving patient outcomes, increasing accessibility to healthcare resources, and enhancing overall patient satisfaction and well-being. Numerous datasets and benchmarks, such as MedQA [4], emrQA [6], and MedMCQA [7], have been developed to train and evaluate LLMs in their ability to answer patient questions effectively.
Recent advancements demonstrate LLMs can be organized into chains to tackle more complex interactions [5], for instance, interacting with patients via natural language while utilizing components that enhance the capabilities of LLMs by integrating additional tools and knowledge bases, such as Retrieval-Augmented Generation (RAG) [8] and ReAct [9]. This structure allows the LLM to not only generate answers to questions but also retrieve relevant information, apply reasoning, and perform specific tasks more effectively.
In real-world LLM healthcare applications, the capability to answer patient questions is just one facet of their functionality [10]. Effective interaction requires LLM healthcare chains to actively engage patients by asking relevant and context-aware questions to gather comprehensive information. This includes symptoms, medical history, lifestyle factors, and other pertinent details that are critical for accurate diagnosis and personalized treatment. By actively soliciting detailed patient information, LLM healthcare chains can enrich the context in which they operate, leveraging integrated tools and smart agents to enhance diagnostic accuracy, provide better interventions, and offer more tailored health advice. Despite the critical importance of this interactive capability, the evaluation of LLM healthcare chains’ questioning abilities—specifically, their effectiveness in asking pertinent questions to elicit essential information from patients—remains significantly under-explored [11, 12].
Addressing this gap, our work proposes a novel evaluation framework to assess and benchmark the questioning capabilities of LLM healthcare chains during conversational interactions with patients. HealthQ not only evaluates the quality of the questions posed by the LLM healthcare chains but also examines whether these questions help in obtaining better answers. We employ mutual information validation to demonstrate the relationship between the quality of the questions and the improvement in the answers received. This approach ensures that the LLM healthcare chains can effectively interact with patients, leading to more accurate diagnoses and better healthcare outcomes.
In summary, our contributions are threefold:
  1. 1.
    We conduct the first comprehensive study on the questioning capabilities of LLM healthcare chains, filling a significant gap in healthcare AI research.
  2. 2.
    We introduce a novel dataset generation pipeline, specifically designed to evaluate the question-asking performance of LLM healthcare chains in realistic patient interaction scenarios.
  3. 3.
    We develop and rigorously validate an evaluation framework that integrates multiple LLM chain types and traditional Natural Language Processing (NLP) metrics. Our framework is validated through two key research questions, ensuring its robustness and providing a solid benchmark for future studies on LLM question-asking in healthcare.
We aim for this framework to serve as a platform for researchers to develop general chain models capable of solving complex real-world questions through extensive logical reasoning in both retrieval and comprehension stages. This approach will enhance the effectiveness and reliability of LLM healthcare chains in diverse healthcare scenarios. To further contribute to this area, we will publish our code and datasets.
2 Related Work
2.1 Digital Healthcare
The advent of digital healthcare interventions has revolutionized patient care by leveraging technologies such as wearable devices [13, 14, 15], AI-driven analytics [5, 16], and advanced decision-making frameworks for healthcare resource allocation and optimization [17, 18, 19, 20]. These interventions enable remote monitoring, personalized treatment plans, and predictive analytics, improving patient outcomes and reducing healthcare costs [21, 22]. AI integration, including machine learning and natural language processing for diagnostic accuracy and patient management [23, 1], as well as parameter-efficient fine-tuning approaches [24], are further enhancing healthcare efficiency. Moreover, modeling techniques and representation learning have shown promise in optimizing healthcare systems [3], emphasizing the transformative potential of digital healthcare.
2.2 LLMs in Healthcare
LLMs such as GPT-4 have shown remarkable proficiency in generating human-like text and understanding complex language constructs [25]. These capabilities have been harnessed in various healthcare applications, including automated medical record summarization, virtual health assistants, and clinical decision support systems [26, 27]. Automated medical record summarization condenses patient records into concise summaries, saving time for healthcare providers. Virtual health assistants interact with patients using natural language, providing answers to medical queries and reminders for medication, thereby enhancing patient engagement. Clinical decision support systems leverage the extensive knowledge embedded in LLMs to assist in diagnosing diseases and suggesting treatment options. LLMs streamline clinical workflows and improve patient engagement by providing accurate and timely responses to medical queries, ultimately enhancing the overall quality of care.
2.3 Interactive Questioning with LLM Healthcare Chains
Interactive healthcare conversational systems are crucial in providing patients with accurate and relevant information. Datasets and benchmarks like MedQA [28] and MedMCQA [7] have played a significant role in advancing QA systems within the medical domain, focusing on optimizing the accuracy and relevance of answers generated by LLMs. These systems use large datasets to fine-tune model responses, making them contextually relevant and patient-centered. However, while answer generation has been prioritized, strategies for optimizing question formulation remain largely unexplored.
Several approaches have been introduced to refine the questioning capabilities of LLMs. Techniques such as Chain of Thought (CoT) [29] prompting, reflective chains [30], and Retrieval-Augmented Generation (RAG) [8] integrate reasoning processes that help generate more contextually relevant questions. CoT prompting allows LLMs to break down complex queries into intermediate steps, leading to more precise and targeted questioning. Reflective chains enable the model to iteratively improve the questions by reflecting on previous intermediate answers, thus enhancing the relevance of subsequent queries. RAG, on the other hand, combines retrieval and generation processes to ensure that questions are informed by the most relevant information retrieved from a knowledge base. Despite these advancements, most existing work has primarily focused on optimizing task completion and response accuracy, with less attention given to the development of sophisticated questioning strategies that can drive more informative patient interactions. Our work addresses this gap by specifically evaluating the questioning capabilities of LLM healthcare chains, contributing a novel framework that benchmarks their ability to engage in meaningful and effective patient dialogues.
3 Method
3.1 Overview of HealthQ
For benchmarking, we customized several state-of-the-art LLM healthcare chains to simulate a doctor’s role, using training data for search and retrieval without prior knowledge of test patients. Virtual patients were simulated using test data, and the LLM chains’ generated questions were assessed. The evaluation framework includes generating patient statements, scoring questions, and comparing results with traditional NLP metrics. HealthQ uses two primary evaluation metrics: LLM judge interrogation and summarization-based evaluation. LLM judge interrogation involves an additional LLM evaluating question quality based on relevance, specificity, and informativeness. Summarization-based evaluation assesses how well-generated answers capture the complete patient case by comparing them to detailed patient information from the test data.
Figure 1 illustrates our system framework. Original medical notes are split into train and test sets. The train set builds a vectorized knowledge base. The test set, parsed by an LLM parser, produces ground truth pairs. The LLM healthcare chain generates questions using additional tools. A virtual patient simulator provides partial patient statements and answers. The generated questions are evaluated by an LLM judge based on various metrics. Finally, answers are assessed using summary-based metrics like NER2NER and ROUGE.
We also analyze the relationship between question quality and answer comprehensiveness through correlation and mutual information studies, demonstrating how well-formulated questions enhance patient information gathering and diagnostic accuracy.
3.2 Data
HealthQ necessitates a general data processing mechanism due to the unstructured nature of medical notes and the often ambiguous doctor-patient conversations. Patients may provide simple ”yes” or ”no” answers, leaving the context or symptoms embedded in the doctor’s questions. To accurately evaluate the questioning capabilities of LLM healthcare chains, we require a customized data format that aligns with our evaluation framework. This involves clearly separating patient known knowledge—which includes the patient’s feelings, self-reported symptoms, and background history—from the doctor’s question statements. These pairs are essential for our virtual patient mechanism, enabling the assessment of how effectively doctors ask questions to gather useful information for diagnosis. To illustrate how the dataset format changed through our processing, Table 1 shows an example of the original and processed data formats.
3.2.1 Dataset
To evaluate the questioning capabilities of LLM healthcare chains, we utilized two public datasets: ChatDoctor [31] and MTS-Dialog [32]. The ChatDoctor dataset comprises 110,000 anonymized and grammatically corrected medical conversations, covering a wide range of symptoms, diagnoses, and treatment plans. This dataset, originally used to fine-tune a medical chat model based on LLaMA [33], provides a robust foundation for assessing the LLM chains’ ability to generate relevant and context-aware questions. The MTS-Dialog dataset includes 1,700 doctor-patient dialogues and corresponding clinical notes, capturing detailed exchanges across various medical specialties. These clinical notes summarize patient encounters, highlighting key medical facts, diagnoses, and treatment plans. The comprehensive nature of this dataset allows for a thorough assessment of the LLM chains’ questioning capabilities in diverse medical scenarios. By utilizing these two datasets, we ensure a diverse and representative evaluation, covering a broad spectrum of medical conditions and conversational contexts.
Table 1: Example of original and processed data. The original format includes detailed patient background, symptoms, diagnosis, history, plan of action, and dialogue. The processed format extracts essential patient known knowledge and reformulates doctor question statements for clarity and consistency.
Original Data Format
Background:
Symptoms: Congestion, increased coughing, fever (temperature of 101°F).
Diagnosis: Upper respiratory infection, possible pertussis.
History of Patient: Patient experiencing congestion, increased coughing, and fever; possibility of pertussis indicated by doctor, inquiring about presence of apnea.
Plan of Action: Doctor to evaluate and determine appropriate course of action.
Dialogue:
Doctor: How’s he feeling today?
Guest_family: I think this is the worst he’s been feeling all week.
Doctor: I’m sorry to hear that he hasn’t been feeling well. What symptoms has he been having?
Guest_family: He’s been very congested as of late and seems to be coughing a lot more than usual. He was also running a fever yesterday.
Doctor: I see. What was his temperature?
Guest_family: The thermometer was reading one hundred and one degrees Fahrenheit. Does he need to go to the hospital?
Doctor: Let me evaluate and see what we can do for him today.
Guest_family: Thank you, doctor.
Doctor: I suspect that he has an upper respiratory infection, possible pertussis. Is he still experiencing apnea?
Processed Data Format
Patient Known Knowledge:
The patient has been very congested lately and coughing more than usual. He had a fever of 101°F yesterday.
Doctor Question Statements:
The doctor asked about the patient’s symptoms and temperature. The doctor suspects the patient has an upper respiratory infection, possibly pertussis, and asked if the patient is still experiencing apnea. The doctor said he will evaluate and see what they can do for the patient today.
3.2.2 Data Processing
As depicted in Figure 1, we split the original medical notes into training and test sets. The training set is transformed and stored in a vector database (VDB), which the LLM Chain can use for searching and retrieval. However, the LLM Chain has no knowledge of any new patients in the test set. The test set is parsed by the LLM to simulate virtual patients, allowing us to evaluate how effectively these LLM healthcare chains, acting as doctors, can ask relevant questions and make judgments based on the simulated patient responses.
In more specific, each training dataset is processed to create a VDB for searching, implemented using FAISS [^4] https://github.com/facebookresearch/faiss. The customized collate functions can specify comparison fields and utilize metadata for enhanced information retrieval. This flexibility accommodates diverse datasets with varying forms and topics. During agent executions, multiple VDB searching functions can be provided as tools without merging, allowing the LLM to select the appropriate database for searching, also allowing for heterogeneous search functions. VDBs work by finding the nearest vectors, which are represented by embedding models. Each vector database entry is defined as:
$V_{i}=[\text{Record}_{i},\mathbf{v}_{i}],$
where $\mathbf{v}_{i}$ is the vector representation used for comparison to find the nearest vectors in the VDB for efficient medical knowledge retrieval.
Each entry in the vector database, referred to as a record, is structured to include both the content and the associated metadata. The collate function processes the training data to create these records:
$\text{Record}_{i}=(\text{Content}_{i},\text{Metadata}_{i})=\text{Collate}(\text{Data}_{\text{train}}(i))$
These databases were critical in facilitating efficient retrieval and processing of medical information, ensuring that the LLM health chains had access to relevant and structured data during their training and evaluation phases. To generate ground truth data from the test set, we employed the Claude [^5] https://claude.ai/ to parse patient notes into structured formats. The parsing function processes the input patient data to produce an output pair consisting of patient known knowledge and doctor question statements.
The final ground truth for testing, $G$, is thus composed of the results from this processing pipeline:
$GT=\{\text{LLM\_Parser}(\text{Data}_{\text{test}}(i))\}$
3.3 Baseline LLM Healthcare Chains
HealthQ includes several state-of-the-art chains, each with unique mechanisms for generating diagnostic questions. Below are the detailed descriptions of each chain implementation, followed by a combined pseudo-code representation.
Hardcoded Workflow: This is the lightest chain that does not require LLM inference but only uses LLM embedding and NER in a pre-defined workflow. Initially, it takes in a patient’s first statement and retrieves relevant cases from the vector database using a search function. NER is applied to extract symptoms and medications from the retrieved cases. It then confirms each extracted symptom and medication with the patient using predefined templates and updates the state accordingly.
RAG  [8]: This chain integrates retrieval and generation processes. It uses a retrieval mechanism to gather relevant cases from the vector database and then employs a language model to generate questions. The questions are designed to be concise and specific, leveraging the retrieved information to enhance their relevance and utility. We compare two approaches: one using an existing framework, LangChain RetrievalQA [^6] https://docs.smith.langchain.com/old/cookbook/hub-examples/retrieval-qachain, and another customized version tailored specifically for generating questions rather than answers, aligning with the goals of our evaluation framework.
RAG with Reflection [30]: Extending the basic RAG framework, this chain includes a reflection mechanism. After generating an initial question, the chain reflects on the quality and relevance of the question by re-evaluating the context of the retrieved cases. This reflective process allows the chain to refine and improve the question, making it more targeted and contextually appropriate.
RAG with CoT [29]: This chain incorporates a CoT approach, involving iterative reasoning steps. It generates an initial question(or intermediate thinkings) and then performs several iterations of reasoning to refine the question. Each iteration considers additional contextual information from the retrieved cases, progressively enhancing the question’s specificity and relevance.
RAG with Reflection and CoT-Self-Consistency (CoT-SC) [34]: Building on the CoT approach, this chain introduces self-consistency checks. It branches several CoT from the same query, and gather the results of which to see of the different branches reaches similar results. This ensures that the final question is not only refined but also consistent and robust, reducing the likelihood of contradictions. Different from the original CoT-SC which uses a voting approach for the final result, our version uses LLM to compare the several branches of the final results of CoTs to cater for phrasing text output.
ReAct [9]: This chain leverages an LLM prompted with tasks and tools, where the tools are functions described for LLM to call. For example, searching among the training set can be a tool. In each step, the LLM decides whether it can reach an answer or needs to use a tool. If a tool is needed, the LLM specifies the tool name and argument for an executor to call. The executor runs the corresponding functions and feeds the output back into the LLM input for the next step. To highlight the importance of the agent’s ability to choose from more than one tool, we included our VDB search function as one of the tools and added the medical abbreviation lookup tool, medialpy [^7] https://pypi.org/project/medialpy/, as another. We used LangChain’s default implementation to facilitate this process. The ReAct chain’s capability to use heterogeneous sources makes it unique among the chains compared in this study.
Table 2: LLM Judge Interrogation Metrics (Scores range from 0 to 10)
Metrics Description
Specificity Measures how specifically the question targets the patient’s symptoms and conditions.
Usefulness Assesses the utility of the question in providing relevant information for diagnosis.
Relevance Evaluates the pertinence of the question to the patient’s symptoms.
Coverage Determines the extent to which the question covers the information available in the ground truth data.
Fluency Rates how well the question is formulated in fluent, natural language.
3.4 Evaluation Framework
Our evaluation framework employs three primary modules:
  • 1.
    Virtual Patient Simulation: This module simulates the patient’s first statement as the input to the LLM healthcare chains, generating responses based on the first generated question. The simulation mimics real-world patient interactions to provide valid input for the LLM healthcare chains. Specifically, $VPS$ represents the virtual patient simulator function that, given the patient knowledge $K$, generates the initial statement $S_{init}$ (i.e., $VPS:(.,K)\rightarrow Stmt$). Similarly, $VPS$ is used to generate a simulated answer from the LLM doctor’s question $Q$ based on the patient knowledge $K$ (i.e., $VPS:(Q,K)\rightarrow Ansr$). The topic and factual content of the virtual patient are controlled by the patient knowledge $K$, ensuring that the distribution of topics matches the test set. We prompt the virtual patient to generate only short conversations, providing information solely based on the given knowledge $K$. The source of bias will be the LLM phrasing and inconsistency from $K$.
  • 2.
    LLM Judge Interrogation: This metric evaluates the questions generated by the LLM healthcare chains based on specificity, usefulness, relevance, coverage, and fluency, as shown in Table 2. An LLM judge $J$ assesses each question $Q$ using the patient knowledge $K$, providing scores according to these criteria. The interrogation metrics are calculated as follows:
    Interrogation $\text{Interrogation Metrics}=J_{K}(Q_{i,j},K_{i})$
    where $i$ is the index for the patient case and $j$ is the index for the specific question generated for that patient case.
  • 3.
    Summarization-Based Evaluation: This metric evaluates the quality of information retrieved by the generated questions using techniques including Recall-Oriented Understudy for Gisting Evaluation (ROUGE) and Named Entity Recognition (NER)-based similarity. These tools assess how well the simulated patient answers encapsulate the entire patient case by comparing the generated responses to the comprehensive patient information from the test data.
Figure 1: System framework for evaluating the questioning capabilities of LLM healthcare chains.
The evaluation process involves the following steps:
  1. 1.
    Initialize Ground Truth Data ($GT$): Collect comprehensive information about the patient’s symptoms and background to create a dataset for evaluation.
  2. 2.
    Provide Patient Known Knowledge ($K$): Share a subset of the ground truth $GT$ with the LLM healthcare chain, containing the partial knowledge the patient initially provides to the doctor.
  3. 3.
    Provide a Virtual Patient Statement ($Stmt=VPS(.,K)$): Simulate the initial statement a patient would provide to the doctor using an LLM to generate this statement based on the partial knowledge, as patients typically do not provide all information at once. We retain the complete knowledge for evaluation purposes in later steps.
  4. 4.
    Generate Question ($Q$): The LLM healthcare chain generates question $Q$ based on the virtual patient’s first condition statement ($Stmt$).
  5. 5.
    Evaluate the Question ($Q$) using Judge ($J(Q,K)$): Evaluate the generated question using an LLM judge based on criteria such as relevance, specificity, and informativeness, as detailed in Table 2.
  6. 6.
    Generate Simulated Answer and Evaluate as Summary ($Ansr:=VPS(Q,K)$): The LLM virtual patient generates an answer based on the initial question and the known patient information. We then evaluate how well this simulated answer encapsulates the patient’s entire known knowledge ($K$) to determine the quality and completeness of the response.
In this process, a simulated patient generates an initial statement ($Stmt$), which the LLM healthcare chain uses to propose questions. The evaluation system assesses these questions based on specific criteria, with access to additional ground truth data that the healthcare chains do not have. Although the healthcare chains can use tools and reflective processes, they are only given the $Stmt$, and no other test data is provided. The tools are derived from training sets or other external sources, not from the test ground truth. The simulated input ($K|Stmt$), which contains less information than the full ground truth ($GT$), is shared. This ensures that the judges have complete knowledge of the test case, while the healthcare chains work within a fair and restricted evaluation process.
Table 3: Comparison of Summary-Based Metrics across LLM Chains
(a) Information Extraction Quality - NER2NER
LLM Healthcare Chain Medication (%) Symptoms (%) Total (%)
Hardcoded 0.0022 0.0384 0.0317
RAG_default 0.0167 0.0942 0.0727
RAG 0.0126 0.0836 0.0777
RAG_reflection 0.0426 0.1434 0.1350
RAG_reflection_CoT 0.0222 0.1054 0.0960
RAG_reflection_CoT_SC 0.0280 0.0825 0.0720
ReAct 0.0000 0.0072 0.0066
(b) Summarization Performance - ROUGE
LLM Healthcare Chain rouge-1 F1 rouge-2 Recall rouge-2 Precision rouge-2 F1 rouge-L Recall rouge-L Precision rouge-L F1
Hardcoded 0.0382 0.0000 0.0000 0.0000 0.0280 0.0647 0.0369
RAG_default 0.1426 0.0241 0.0156 0.0164 0.1853 0.1119 0.1262
RAG 0.0991 0.0089 0.0170 0.0106 0.0716 0.1290 0.0861
RAG_reflection 0.1931 0.0652 0.0459 0.0480 0.2077 0.1649 0.1675
RAG_reflection_CoT 0.1447 0.0272 0.0245 0.0216 0.1337 0.1430 0.1230
RAG_reflection_CoT_SC 0.1275 0.0170 0.0130 0.0128 0.1243 0.1125 0.1064
ReAct 0.0840 0.0065 0.0049 0.0046 0.0856 0.0840 0.0756
3.5 Validation
To ensure the validity of our evaluation, we devised a framework distinct from traditional interrogation or cross-validation methods. Our approach does not directly rely on ground truth from a test set to evaluate the system. Instead, it employs separate LLM workflows for the roles of judge and doctor. The judge LLM, which has access to the test ground truth data, is responsible for evaluating the questions generated by the LLM healthcare chains. In contrast, the LLM healthcare chains, acting as doctor LLM, only have access to the virtual patient initial statement $Stmt$. This separation ensures that the question generation process remains unbiased and prevents data leakage.
To maintain the integrity of the evaluation, we strictly prevent data leakage. The LLM healthcare chains never directly access the test patient known knowledge $K$. Thus, the phrasing of the questions is based solely on a subset of $K$ and the search results from the chains’ own knowledge base. This design ensures that the evaluation of the generated questions is impartial and based on the information provided in the virtual patient statement.
HealthQ addresses two critical research questions (RQs) to validate its effectiveness:
  • 1.
    RQ1: How well does the induced answer to the question summarize all the relevant information? This question assesses the completeness and quality of the answers generated by the LLM virtual patient in response to their own questions.
  • 2.
    RQ2: How are the results of LLM judge interrogation for the proposed questions correlated with the quality of the induced answers? This question explores the relationship between the quality of the questions and the quality of the answers they elicit, using well-established summarization metrics.
Table 4: LLM Judge Interrogation Results
LLM Healthcare Chain Specificity Usefulness Relevance Coverage Fluency
Hardcoded 3.56 4.52 4.66 2.30 5.53
RAG_default 5.30 5.95 6.50 3.58 8.06
RAG 7.78 7.87 8.75 4.55 9.16
RAG_reflection 7.78 7.94 8.96 5.08 8.92
RAG_reflection_CoT 7.84 8.08 8.98 5.02 9.02
RAG_reflection_CoT_SC 7.93 8.44 9.29 5.34 8.88
ReAct 3.55 4.65 3.46 2.26 7.59
For RQ1, we evaluate the effectiveness of the questions using a virtual patient and established summarization metrics, measuring how well the induced answers encapsulate patient information. For RQ2, we analyze the correlation and mutual information between the quality of the questions and the comprehensiveness of the answers. A strong correlation would suggest that well-formulated questions lead to more informative answers, thereby enhancing patient information gathering and improving healthcare outcomes.
Mutual information $I(X_{i},X_{j})$ quantifies shared information between variables without the need of predictive modeling. By using its standardized form $\frac{I(X_{i},X_{j})}{\min(H(X_{i}),H(X_{j}))}$, where $H(X_{i})$ is the entropy of $X_{i}$ and $H(X_{j})$ is the entropy of $X_{j}$, we identify dependencies or shared information within the interrogation and summarization groups. This ratio indicates the proportion of shared information relative to the smaller entropy value, providing a normalized measure of pairwise dependence. Although discretizing data into bins may introduce numerical errors, the coarsened scores can still effectively achieve the evaluation goal, helping us understand the inherent relationships in our evaluation.
Figure 2: LLM judge interrogation metrics: kernel density estimation and radar plots
Our mutual information validation comprises two steps:
  1. 1.
    Individual Data Point Analysis: This analysis examines the relationship between the interrogation score for each individual question generated by the LLM healthcare chains and the informativeness of the corresponding response from the virtual patient. Specifically, it assesses whether a high score in interrogation metrics (see Table 2) for a single question leads to a more comprehensive and informative answer for that specific interaction.
  2. 2.
    Aggregate Analysis: This analysis evaluates the overall performance of each LLM healthcare chain by averaging the interrogation scores across all interactions. It then investigates whether the average interrogation score for a given LLM healthcare chain correlates with the overall quality and comprehensiveness of the answers induced by that chain across the entire dataset. This helps in understanding whether LLM chains that consistently generate high-quality questions also tend to elicit more informative responses in general.
The results of these analyses, presented through correlation and mutual information matrices in Section 4, provide insights into how well-formulated questions can enhance patient interactions and improve healthcare outcomes.
4 Experiments
4.1 Experiment Setup
To ensure the feasibility of our approach for local implementations in clinics and hospitals, all components of the LLM health chains are based on open-source models. The primary LLM used is the Mixtral [^8] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 8X7b, served on Groq [^9] https://wow.groq.com/. On branches of Chain of Thought Self-Consistency (CoT-SC) processes, a temperature of 0.3 is used, while a temperature of 0 is applied for all other models. The NER component utilizes an open-source biomedical NER model [35] to extract medical entities to make set comparisons. We employ FAISS [36] with its default settings as the VDB index package, and the MiniLM [^10] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 model for text embeddings. Data preprocessing—including the extraction of patient known knowledge ($K$), virtual patient simulation, and question evaluation by the LLM judge—is conducted using Claude-3-Opus [^11] https://claude.ai/. The experiments are conducted on a dataset comprising 128 test cases (64 from each of the two datasets). We benchmark seven different LLM health chains, generating a total of 796 patient status-doctor statement pairs. For comparative analysis, we group the data by the LLM chain type and summarize the mean scores in Table 4.
(a) Individual interactions
(b) Aggregated within specific LLM healthcare chains
Figure 3: Mutual information analysis - correlation between question quality and answer quality, across all metrics
This setup rigorously evaluates the questioning capabilities of various LLM healthcare chains, ensuring comprehensive and reliable results.
4.2 Results and Analysis
This section presents the evaluation results of the LLM healthcare chains, with explicit connections to the research questions RQ1 and RQ2, which were established to validate the effectiveness of our proposed evaluation framework.
The summary-based metrics presented in Table 3(b) are particularly relevant to RQ1, which assesses how well the induced answers summarize all relevant information. The results show that state-of-the-art chains like RAG_reflection and RAG_Reflection_CoT_SC consistently outperform more primitive methods, such as the hardcoded chain, across nearly all evaluated metrics. This is especially evident in the ROUGE and NER2NER metrics, where these advanced chains demonstrate superior performance. For instance, the RAG_reflection_CoT_SC chain achieves the highest overall scores in several categories, reflecting its ability to generate contextually relevant and informative questions that lead to more comprehensive and accurate patient responses. The lower performance of the ReAct chain could be attributed to its design in LangChain, emphasizing answering rather than questioning, along with the limited number of tools provided to it. These findings suggest that well-formulated questions by the more advanced chains lead to better information extraction, thereby validating the effectiveness of our evaluation framework in summarizing all relevant information, as addressed in RQ1.
For RQ2, which examines the correlation between the quality of questions generated by LLM chains and the resulting answers, the LLM judge interrogation results in Table 4 show that most LLM healthcare chains substantially outperform the hardcoded workflow across all interrogation metrics. The basic RAG chain shows notable improvements in specificity, usefulness, relevance, and fluency compared to the hardcoded method, highlighting its more flexible and adaptive approach to generating questions. Chains that incorporate reflection and iterative reasoning techniques, such as those using reflection and CoT, exhibit further enhancements, particularly in relevance and coverage. The reflection mechanism enables the chains to reassess and refine the context of the questions, leading to more targeted and contextually appropriate queries. The addition of self-consistency checks further boosts performance, with the RAG_CoT_SC chain showing the highest scores in relevance and fluency. This improvement is attributed to the iterative reasoning and self-consistency processes, which help in formulating more coherent and relevant questions. The reaction-based chain, while showing some improvements over the hardcoded method, still lags behind the more advanced RAG variants, particularly in specificity and coverage. This analysis demonstrates that chains performing well in interrogation metrics tend to produce better quality answers, thus validating the connection explored in RQ2.
Figure 2 presents the kernel density estimation (KDE) and radar plots comparing the distributions of interrogation metrics across LLM healthcare chains, providing further insights into RQ2. The KDE plots reveal that advanced chains, such as RAG and RAG with reflection, exhibit broader and higher-density distributions in key metrics like specificity, relevance, and usefulness. This reflects their adaptive questioning capabilities and iterative reasoning processes. Chains incorporating reflection and CoT techniques show distinct improvements, particularly in relevance and specificity, due to their ability to refine questions based on previous outputs. The radar plot consolidates these findings, with self-consistency-enhanced chains (RAG_reflection_CoT_SC) displaying more balanced performance across all metrics. In contrast, the hardcoded chain has narrow distributions, indicating rigid question generation, while the ReAct chain struggles with fluency and coverage, possibly due to limitations in its toolset. These observations support the idea that well-designed chains, especially those employing reflection and CoT methods, are more effective in generating questions that lead to informative patient responses, aligning with the goals outlined in RQ2.
Figures 3(a) and 3(b) offer deeper insights into the relationship between question quality and answer quality using normalized mutual information (MI) matrices. In Figure 3(a), MI is computed between two metrics, $J_{m}$ and $J_{n}$, for each individual interaction:
$MI\left(J_{m}(data_{p}),J_{n}(data_{p})\right)$
where $p$ indexes individual interactions or patient cases. This figure shows strong dependence between metrics such as ”Usefulness” and ”Relevance” ($MI=0.50$), reflecting their shared focus on capturing relevant aspects of the patient’s condition. In contrast, pairs with less dependency such as those between ”Fluency” and ”Coverage” (MI = 0.38) indicate weaker relevance. In this case, fluency does not necessarily imply the completeness of aspects covered in the question. In Figure 3(b), MI is calculated based on aggregated scores across all interactions within each LLM chain type:
$MI\left(J_{m}\left(\sum_{i\in Chain[l]}data_{i}\right),J_{n}\left(\sum_{i\in Chain[l]}data_{i}\right)\right)$
Here, the summation $\sum_{i\in Chain[l]}data_{i}$ represents all interactions $i$ for a specific chain type $Chain[l]$. This aggregated analysis reveals that correlations within each chain type remain consistent, with strong relationships among metrics like ”Usefulness” and ”Relevance.” Chains that score well in these areas also tend to perform well in summarization metrics like ROUGE and NER2NER, affirming the robustness of the evaluation framework.
Overall, the results indicate that higher interrogation scores correlate with better summarization performance, confirming our hypothesis that good questions lead to better information extraction. This correlation underscores the importance of effective questioning in improving the overall quality of patient interactions and diagnostic outcomes, thereby demonstrating the value of our evaluation framework.
5 Discussion
5.1 Sources of Bias and Correction
Our evaluation framework considers potential biases in both the sampling process and the construction of $Stmt$ and $GT$.
5.1.1 Pre-judgement Errors
Randomized train-test splits and sampling from both ChatDoctor and MTS-Dialog datasets can lead to distributional biases. Re-weighting patient types can help mitigate this issue. Errors may arise from the LLM’s data cleaning process due to hallucinations, or from patient case data that does not fully capture essential diagnostic information. We assume the patient data is comprehensive and accurately reflects real-world doctor-patient interactions. It may be possible to reduce this bias through conditioning on patient profiling and matching them with simulation promptings. Bias can also occur in transitioning from $K$ to $Stmt$, as patient statements may vary. Addressing this as a sample bias problem with fixed conversation starters can help correct this.
5.1.2 Judgement Errors
Interrogation biases stem from LLM prompting and judgement errors, while summarization metrics biases arise from LLM-simulated answers. Mitigating these biases involves diversifying error sources by using LLM capabilities differently: one for scoring directly and another for extracting $Q$ from $K$. Effective evaluation should align with good diagnosis and ultimately improve patient outcomes.
5.2 Multi-shot Conversation with Memory
Real clinical interactions involve multiple QA rounds, which our current study does not fully capture due to text length and computational limitations. Future work with longer context windows and richer patient data can address this. As conversation complexity increases, it may be modeled as a dynamical system to study more intricate behaviors.
6 Conclusion
This paper presents a novel framework for evaluating the questioning capabilities of LLM healthcare chains, addressing a significant gap in the current research landscape. By implementing various state-of-the-art LLM chains and utilizing traditional NLP metrics for validation, we provide a comprehensive assessment of their ability to generate relevant and informative questions during patient interactions. Our approach includes the development of custom datasets and an evaluation framework that separates patient known knowledge from doctor questions, enabling precise evaluation. The results indicate that well-formulated questions enhance patient information gathering, thereby improving diagnostic accuracy and healthcare outcomes. Future work will focus on expanding the framework to support multi-shot conversations and further reducing biases in evaluation metrics. This framework sets a benchmark for future studies and aims to advance the effectiveness of LLM healthcare chains in real-world applications.